OpenStudy (anonymous):

Assume that the procedure yields a binomial distribution with trial repeated n times. Use the binomial probability formula to find the probability of x successes given the probability p of success on a single trial. n=5, X=2 , p=0.45 Thanks!

OpenStudy (anonymous):

As I said, I know little about probability calculations. Perhaps you should ask @bhaskarbabu, as he claims to have a firm grasp on it.

OpenStudy (anonymous):

I am asking you a favor not him! So don't worry about it!

OpenStudy (anonymous):

Can you help me, Plzzz!!!

OpenStudy (anonymous):

Okay, write out your formula. Determine what n, X, and p represent.

OpenStudy (anonymous):

ok! hold on!

OpenStudy (anonymous):

my compu keep frezing! not sure whats the problem.

OpenStudy (anonymous):

I had to change the question..? I got it wrong and they gave the new.

OpenStudy (anonymous):

That's fine. The equation stays the same.

OpenStudy (anonymous):

Here is what I got! p(3)=(5!)/(5-3)!3!= 0.65^(3)*0.35^(3)

OpenStudy (anonymous):

Then (65)(0.2746)(0.042875) Then I got stuck here.

OpenStudy (anonymous):

Okay. I'm not sure about your use of factorials, but that's not the formula I'm familiar with.

OpenStudy (anonymous):

I know of: \[P = \left(\begin{matrix}n \\ k\end{matrix}\right) p^{k} q^{n-k}\]

OpenStudy (anonymous):

No, it's different, they gave me the formula! you want me to give to you?

OpenStudy (anonymous):

Please.

OpenStudy (anonymous):

P(x)=(n!)/((n-x)!x!)*p^(x)*q^(n-x)

OpenStudy (anonymous):

Let me make sure I understand you. Is it: \[P_{x} = \frac{n!}{((n-x)! x!)} p^{x} q^{n-x}\] or \[P_{x} = \frac{n!}{((n-x)! x!) p^{x} q^{n-x}}\]

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Which one? They're very different.

OpenStudy (anonymous):

the first one..

OpenStudy (anonymous):

Okay. What does q stand for?

OpenStudy (anonymous):

q=1-p

OpenStudy (anonymous):

in this case new question it will be 1- 0.15=0.85

OpenStudy (anonymous):

Am I correct?

OpenStudy (anonymous):

Okay, so you should be able to plug in all your values. Earlier, you wrote p(3)=(5!)/(5-3)!3! but you seem to have forgotten to multiply it by p^x q^(n-x)

OpenStudy (anonymous):

Ok, I figured it out! But how do I multiply.

OpenStudy (anonymous):

The way that I did I got 30..

OpenStudy (anonymous):

p(2)=(5!)/(5-2)!2!

OpenStudy (anonymous):

Go back to your formula. You're missing everything after the fraction. \[P_{x} = \frac{n!}{((n−x)!x!)} p^{x} q^{n−x}\]

OpenStudy (anonymous):

P(2)=(5!)/((5-2)!2!)*0.15^(2)*0.85^(5-2)

OpenStudy (anonymous):

@geoffb Here is what I got. Can you explain me why they the right answer is 0.138?? Thank you!

OpenStudy (anonymous):

Is this how you wrote out your formula on paper? \[P = \frac{5!}{3! \times 2!} (0.15)^{2} (0.85)^{3}\]

OpenStudy (anonymous):

How you calculate the fraction only?

OpenStudy (anonymous):

I got 120/8 is that correct?

OpenStudy (anonymous):

Nope. Your numerator is right, but not the denominator. What is 3!? What is 2!? What is the product of the two?

OpenStudy (anonymous):

Ohh 12

OpenStudy (anonymous):

Good. Now are you able to get the answer?

OpenStudy (anonymous):

Nooo!! I have an idea but nop..

OpenStudy (anonymous):

So what I do is divide the fraction after in this case will be (10).

OpenStudy (anonymous):

Okay, so now you have this, right? : \[P = \frac{120}{12} (0.15)^{2} (0.85)^{3}\]

OpenStudy (anonymous):

(10)(0.0225)(0.7225)

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

No, 0.85^3 is not 0.7225.

OpenStudy (anonymous):

0.614

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok! let me see.

OpenStudy (anonymous):

I got it right now..

OpenStudy (anonymous):

My mistake was leaving the whole fraction without realizing It was missing..I am dummi!! jejeeje!!

OpenStudy (anonymous):

No problem. You seem to understand how the calculation works now. :)

OpenStudy (anonymous):

Yes, Thank you sooo muchhh!!!! :)